Please can you help me with this problem: The sum of the ages of dorothy and dorilyn is 41. In 5 years, dorothy will be twice as old as dorilyn. Find thier ages 3 years ago.

Answer by multiplier(8)


say: x= age of dorothy at present
y= age of dorilyn at present
but sum of their ages is 41
therefore:
x+y=41>>>> this is your equation 1
in 5 years, meaning after 5 years, dorothy will be wice a old as dorilyn
from their present ages after 5 years,
dorothy will be (x+5) years old and dorilyn will be (y+5) years old as well
and dorothy will be twice as old as dorilyn
(x+5), the age of dorothy will be equal to twice the age of dorilyn which
is 2(y+5)
(x+5)=2(y+5) simplifying this equation gives: x-2y-5=0>>>> your equation 2
now you have two equations
x+y=41 and x-2y-5=0, two simple equations with two unknown
further solving, we get x=29 years old(dorothy) & y=12 years old(dorilyn)
so to get there ages three years back, subtract 3 from their present ages
dorothy's ages was: (29-3)=26 and dorilyn's age was(12-3)=9
checking: substitute values of x and y in equations 1 & 2
x+y=41, 29+12=41, and x-2y-5=0, 29-2(12)-5=0

hope you'll like my solution